Calculating CPI, real, and per capita terms.
Year: 2000, 2001, 2002.
Data: 847, 902, 911.
GDP ($B): 562, 607, 619.
Pop (m): 183, 201, 232.
Calculate CPI 2000 for 2002.
1
Expert's answer
2016-04-30T05:11:02-0400
Calculating CPI, real, and per capita terms.
Year: 2000, 2001, 2002.
Data: 847, 902, 911.
GDP ($B): 562, 607, 619.
Pop (m): 183, 201, 232.
CPI(2000) = 100,
CPI(2001) = 607/562*100 = 108,
CPI(2002) = 619/607*100 = 102.